OpenStudy (anonymous):

If f(x) = 3a|4x – 4| – ax, where a is some constant, find f ′(1) with the answer options being: 0 not enough information e 1

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i would say "none of the above" since it is unlikely that this function has a derivative at \(x=1\)

OpenStudy (misty1212):

suppose for example \(a=1\) and you had \[f(x)=3|4x-4|-x\] what would the derivative of that be?

OpenStudy (anonymous):

\[f \prime \left( x \right)~does~\not~exist~if~4x-4=0~or~x=1\]

OpenStudy (michele_laino):

if we write, explicitly, the absolute value, we can rewrite the function as below: \[\Large f\left( x \right) = \left\{ {\begin{array}{*{20}{c}} {12a\left( {x - 1} \right) - ax,\quad x > 1} \\ { - a,\quad x = 1} \\ {12a\left( {1 - x} \right) - ax,\quad x < 1} \end{array}} \right.\] now, in order to see if it exists the function \(f'(1)\), please you have to compute these two limits: \[\Large \begin{gathered} \mathop {\lim }\limits_{x \to 1 - } \frac{{f\left( x \right) - f\left( 1 \right)}}{{x - 1}} = ...? \hfill \\ \hfill \\ \mathop {\lim }\limits_{x \to 1 + } \frac{{f\left( x \right) - f\left( 1 \right)}}{{x - 1}} = ...? \hfill \\ \end{gathered} \] of course, in order to do the substitution, you have to use my defintion above of the function